# Emotional Support Evaluation Framework via Controllable and Diverse Seeker Simulator

### v1

@@ 5.3.1 Experimental Setup
To assess each simulator’s coverage of diverse seeker characteristics, we evaluate the diversity of the generated utterances. Using 300 held-out Reddit-derived profiles per simulator, we conduct dialogues with GPT-5-mini as the supporter, following the baselines in Section 5.1.1. Details on simulator-specific profile construction are provided in Appendix F. For each dialogue, seeker utterances are aggregated into a single embedding using the all-MiniLM-L6-v2 to capture the seeker’s overall expressive pattern Reimers and Gurevych (2019).
@@ 5.3.3 Results and Analysis
Figure 3 visualizes the semantic diversity of seeker utterances using UMAP projections. Our simulator exhibits substantially broader coverage compared to all baseline simulators, indicating a wider range of expressed behaviors and interaction styles. In contrast, baseline simulators tend to concentrate within narrower regions of the embedding space, suggesting limited behavioral diversity.
@@ 6 Supporter Model Evaluation
In this section, we utilize our controllable seeker simulator to evaluate the performance of state-of-the-art supporter models, examining how their performance and supportive behaviors vary across diverse seeker populations. To examine the impact of seeker populations on evaluation outcomes, we evaluate the same support models using other existing seeker simulators under an identical settings.
@@ 6.3 Supporter Models
  • Fine-tuned: Llama-3-8B-Instruct models fine-tuned on emotional support datasets, i.e., Llama-ESConv, Llama-ExTES, Llama-Psych8k, Llama-PsyInsight, and Llama-CounselChat (see Appendix M for dataset sources).
@@ 6.4 Evaluation Results
First, supporter model rankings vary substantially across seeker simulators. With cooperative simulators such as ESC-Judge and ESC-Role, models achieve uniformly high ES skills with small performance gaps, whereas under the more challenging seeker behaviors induced by our simulator, ES scores drop sharply and model rankings change noticeably. This suggests that evaluations based on limited seeker populations can overestimate supporter robustness. Second, evaluation metrics differ in their sensitivity to seeker behaviors. While general skills remain relatively stable across simulators, ES skills—particularly suggestions and informativeness—exhibit larger performance drops and higher variance under our simulator. This indicates that challenging seeker behaviors primarily affect a model’s ability to deliver substantive support rather than surface-level language quality. (Table 13)
@@ 6.5 Guidelines for Seeker Profile Configuration
Developers of emotional support chatbots can configure input seeker profiles according to the target seeker population of the chatbot. In Appendix L, we provide guidance on how to configure the seeker profiles depending on whether the chatbot targets general online users or formal clinical psychotherapy settings.
@@ Appendix D Human Validation for LLM-based Tagging
Feature-level inter-annotator agreement (IAA) and human–LLM alignment results are reported in Table 2
@@ E.1 Linguistic Feature Extraction Rules
Appendix F Profile construction for each Seeker Simulator
To ensure fair comparison and preserve the originally reported performance of each seeker simulator, we adopt the original profile structures and prompting schemes proposed in their respective papers without modification. In addition, when constructing simulator-specific profiles, we aim to cover a broad range of seeker characteristics within each simulator’s representational capacity by assigning diverse feature labels where applicable.
ESC-Eval
Eeyore
ESC-Judge
Appendix G Implementation Details for Contrastive Learning
@@ H.2 Expert Evaluation Settings
Appendix I Routing Analysis in MoE
Figure 5 visualizes the routing space via PCA. Distinct clustering by dominant expert suggests that the routing network learns separable regions in $\boldsymbol{\alpha}$ space, reinforcing the observed feature-conditioned specialization. Figure 6 presents conditional distributions of dominant experts given each behavioral feature. These patterns support the claim that experts capture distinct behavioral regimes rather than surface-level variations.
@@ Appendix I Routing Analysis in MoE
Seeker Simulator TSNE Hull Area PCA Hull Area
@@
Appendix L Feature Distribution Analysis on Real-world Emotional Support Context
To illustrate how our framework can be applied in practice, we apply the same feature annotation pipeline to a real-world spoken counseling dialogue dataset Alexander Street (2020). We compare feature distributions between this target seeker population and our Reddit-based training data to identify population-level differences (Figure 10).
These differences reflect variations in the seeker populations targeted by supporter models, offering a concrete basis for deriving target seeker profiles and configuring the simulator for population-aware evaluation.
Appendix M Datasets Used for Fine-tuning Supporter Models
The supporter models were fine-tuned on publicly available emotional support datasets, including ESConv Liu et al. (2021), ExTES Zheng et al. (2024), Psych8k Liu et al. (2023), and PsyInsight Chen et al. (2025), CounselChat Bertagnolli (2020).
@@
Model Coping Utterance Profanity Engagement Verbosity Self-Disclosure Resistance Turns
GPT-4.1-mini 0.33 0.36 0.58 0.27 0.28 0.19 0.24 0.16
Llama-3-8B-Instruct 0.29 0.24 0.58 0.27 0.11 0.14 0.29 0.15
Qwen-2.5-14B-Instruct 0.25 0.32 0.47 0.29 0.33 0.22 0.24 0.15
SFT 0.38 0.36 0.66 0.53 0.66 0.40 0.37 0.56
Contrastive Learning 0.30 0.32 0.64 0.39 0.54 0.32 0.35 0.83
Ours 0.44 0.47 0.63 0.58 0.69 0.45 0.43 0.74
@@
Llama-3-8B-Instruct Eeyore 4.301 4.816 3.870 1.977 3.836 5.000 5.000 4.993 4.993 4.993
@@
Llama-CounselChat Eeyore 3.893 3.850 3.820 1.830 4.060 4.803 4.780 4.530 4.313 4.503
@@
Llama-ESConv Eeyore 3.817 3.927 3.543 2.053 3.133 4.907 4.890 4.587 4.457 4.593
@@
Llama-ExTES Eeyore 4.333 4.660 3.720 1.903 3.797 5.000 5.000 4.990 4.980 4.990
@@
Llama-Psych8k Eeyore 4.097 4.183 3.823 1.850 3.843 4.977 4.983 4.923 4.893 4.927
@@
Llama-PsyInsight Eeyore 3.837 3.563 3.040 1.433 2.537 4.870 4.790 4.653 4.457 4.623

### v2

@@ 5.3.1 Experimental Setup
To assess each simulator’s coverage of diverse seeker characteristics, we evaluate the diversity of the generated utterances. Using 300 held-out profiles per simulator, we conduct dialogues with GPT-5-mini as the supporter. We compare our simulator against four baselines: ClientCAST Wang et al. (2024a), ESC-Judge, ESC-Role, and Eeyore. Details on simulator-specific profile construction are in Appendix I. For each dialogue, seeker utterances are aggregated into a single embedding using the all-MiniLM-L6-v2 to capture the seeker’s overall expressive pattern Reimers and Gurevych (2019).
@@ 5.3.3 Results and Analysis
Figure 3 visualizes the semantic diversity of seeker utterances using UMAP projections. Our simulator exhibits substantially broader coverage compared to all baseline simulators, indicating a wider range of expressed behaviors and interaction styles. In contrast, baseline simulators tend to concentrate within narrower regions of the embedding space, suggesting limited behavioral diversity. Notably, simulators without fine-tuning (ESC-Judge and ClientCAST) occupy the most constrained regions, while SFT-based simulators (ESC-Role and Eeyore) achieve moderately broader coverage. Our MoE-based simulator achieves the largest coverage area, demonstrating that assigning dedicated experts to distinct seeker profiles effectively expands behavioral diversity.
@@ 6 Supporter Model Evaluation
In this section, we utilize our controllable seeker simulator to evaluate the performance of state-of-the-art supporter models, examining how their performance and supportive behaviors vary across diverse seeker populations. To examine the impact of seeker populations on evaluation outcomes, we evaluate the same support models using other existing seeker simulators under identical settings.
@@ 6.3 Supporter Models
  • Fine-tuned: Llama-3-8B-Instruct models fine-tuned on emotional support datasets, i.e., Llama-ESConv, Llama-ExTES, Llama-Psych8k, Llama-PsyInsight, and Llama-CounselChat (see Appendix L for dataset sources).
@@ 6.4 Evaluation Results
First, supporter model rankings vary substantially across seeker simulators. With cooperative simulators such as ESC-Judge and ESC-Role, models achieve uniformly high ES skills with small gaps, whereas under our simulator’s more challenging seeker behaviors, ES scores drop sharply and model rankings change noticeably. This suggests that evaluations based on limited seeker populations can overestimate supporter robustness.
Second, evaluation metrics differ in their sensitivity to seeker behaviors. While general fluency remains relatively stable across simulators, ES skills—particularly suggestions and informativeness—exhibit larger performance drops and higher variance under our simulator. This indicates that challenging seeker behaviors primarily affect a model’s ability to deliver substantive support rather than surface-level language quality (Table 18).
Third, qualitative analysis of low-scoring dialogues reveals two recurring failure patterns. With resistant seekers, models fall back on repetitive apologies instead of providing substantive empathy or constructive suggestions. With low-engagement seekers, they persist in monotonous probing rather than adapting strategies like validation or reflection. These patterns explain the observed drops in Suggestions and Informativeness.
6.5 Validity of Automatic Evaluation
To verify that the observed score drops across seeker simulators reflect genuine deficiencies in the supporter model’s skills, rather than the automated judge assigning low scores in response to the seeker’s negative or resistant behaviors, we conduct a human evaluation. Two graduate students in clinical psychology evaluated 60 dialogues using the same 1–5 rating scale and rubrics as the automated evaluation (details in Appendix M).
Specifically, we compute $\Delta_{\text{Human}}$ and $\Delta_{\text{LLM}}$ as the mean score differences (Ours - ESC-Judge) under human and automated evaluation, respectively. A negative $\Delta$ indicates that supporter models perform worse when evaluated with our simulator compared to ESC-Judge. As shown in Table 3, both $\Delta_{\text{Human}}$ and $\Delta_{\text{LLM}}$ are consistently negative across all five ES-skills, confirming that the score drops are not an artifact of automated scoring. Notably, with the exception of Experience Sharing, $\Delta_{\text{Human}}$ is larger in magnitude than $\Delta_{\text{LLM}}$ across all metrics, indicating that human experts perceive the performance gap even more strongly than the automated judge. This suggests that the score drops observed under our simulator reflect genuine supporter performance degradation, rather than systematic bias introduced by the seeker’s negative behaviors.
Metric $\boldsymbol{\Delta_{\text{Human}}}$ $\boldsymbol{\Delta_{\text{LLM}}}$ Spearman $\rho$
Identification -1.167 -0.733 0.454
Comforting -0.900 -0.833 0.377
Suggestions -1.583 -1.067 0.744
Exp. Sharing -0.417 -1.067 0.402
Informativeness -1.917 -1.467 0.673
Table 3: Score differences ($\Delta$ = Ours - ESC-Judge) under human and automated evaluation, respectively, and their rank correlations (Spearman $\rho$) across ES-skills. All correlations are significant ($p<0.05$).
@@ 6.6 Guidelines for Seeker Profile Configuration
Developers of emotional support chatbots can configure input seeker profiles according to the target seeker population of the chatbot. To illustrate how our framework can be applied in practice, we apply the same feature annotation pipeline to a real-world spoken counseling dialogue dataset Alexander Street (2020) and compare feature distributions between this target seeker population and our Reddit-based training data (Figure 10). These population-level differences offer a concrete basis for configuring target seeker profiles. For instance, a chatbot targeting general online users may configure profiles based on Reddit-based interactions, where seekers tend to communicate in a plain and concise style with varying levels of self-disclosure. In contrast, a chatbot targeting formal clinical psychotherapy settings may prioritize verbose utterance styles and lower self-disclosure levels (levels 1–2), reflecting the more structured and surface-level communication patterns observed in spoken counseling interactions.
@@ 7 Conclusion
Acknowledgments
This work was supported by the Creative-Pioneering Researchers Program through Seoul National University and by the National Research Foundation of Korea (NRF) grants (RS-2024-00333484 and RS-2024-00414981) funded by the Korean government (MSIT).
We used AI assistants to proofread the writing and to help with coding.
@@ Appendix D Human Validation for LLM-based Tagging
Feature-level inter-annotator agreement (IAA) and human–LLM alignment results are reported in Table 4.
@@ E.1 Linguistic Feature Extraction Rules
Appendix F Implementation Details for Contrastive Learning
@@
Appendix G Routing Analysis in MoE
Figure 5 visualizes the routing space via PCA. Each point corresponds to a dialogue sample and is colored by its dominant expert $e^{*}=\arg\max\boldsymbol{\alpha}$. Distinct clustering by dominant expert suggests that the routing network learns separable regions in $\boldsymbol{\alpha}$ space, reinforcing the observed feature-conditioned specialization. This structural separation indicates that routing decisions are organized in a meaningful latent space, where each expert occupies a consistent region, supporting expert-level interpretation. Figure 6 presents conditional distributions $P(\text{feature}\mid e^{*})$ of behavioral features given the dominant expert. Each row is normalized to sum to 1. High values in a row indicate that a particular feature level is strongly associated with that expert. These patterns support the claim that experts capture distinct behavioral regimes rather than surface-level variations.
These distributions reveal distinct behavioral tendencies for each expert. For example, Expert 1 shows strong concentration on high engagement (1.0), low resistance (0.76), and high self-disclosure (level 4: 0.53), whereas Expert 3 exhibits relatively higher proportions of low engagement and lower self-disclosure levels (1–2). Such patterns demonstrate that expert routing aligns with coherent behavioral profiles rather than superficial variation.
Despite this expert-based structure, the MoE framework incurs minimal overhead. The routing network adds only 15,881 parameters ($\sim$}0.0003% of total model size), requiring approximately 2 hours and 40 minutes of additional training on a single H100 GPU. Moreover, MoE provides targeted improvements on features that SFT largely fails to control, such as self-disclosure level (0.40 $\rightarrow$ 0.45) and resistance level (0.37 $\rightarrow$ 0.43), effectively expanding the range of reliably controllable behavioral dimensions.
@@ H.2 Expert Evaluation Settings
Annotator Details
The evaluation was conducted by three graduate students specializing in clinical psychology (two males and one female). Participation was voluntary, and each annotator was compensated at approximately \$40 per hour in accordance with standard research participation guidelines.
Inter-Annotator Agreement
We report inter-annotator agreement to assess the reliability of our expert evaluations. Annotators were asked to choose among “prefer A”, “prefer B”, and “prefer both”. Given the ordinal nature of the scale (A = 0, Both = 0.5, B = 1), we compute mean absolute ordinal deviation to capture partial agreement rather than treating all disagreements as equally distant. The resulting scores were 0.341, 0.315, and 0.322 for Linguistic Naturalness, Role Authenticity, and Psychological Plausibility, respectively, indicating that annotators rarely showed completely opposite preferences. Pairwise agreement and full agreement rates are reported in Table 6.
Metric 1–2 1–3 2–3 All
Linguistic Naturalness 0.811 0.433 0.411 0.367
Role Authenticity 0.744 0.411 0.389 0.311
Psychological Plausibility 0.800 0.300 0.311 0.267
Table 6: Pairwise and full inter-annotator agreement rates. Columns 1–2, 1–3, 2–3 denote pairwise agreement between annotators, and All denotes the rate of full agreement across all three annotators.
Appendix I Profile construction for each Seeker Simulator
To ensure fair comparison and preserve the originally reported performance of each seeker simulator, we adopt the original profile structures and prompting schemes proposed in their respective papers without modification. We construct 300 held-out Reddit-derived profiles for each simulator following its originally proposed profile construction procedure. For ClientCAST, we instead use its originally proposed datasets directly, as the simulator requires reference dialogues to be provided during generation. In addition, when constructing simulator-specific profiles, we aim to cover a broad range of seeker characteristics within each simulator’s representational capacity by assigning diverse feature labels where applicable.
ESC-Judge
ESC-Eval
Eeyore
ClientCAST
For ClientCAST, we reproduce the simulator following the methodology and codebase described in the original paper. Seeker profiles are constructed from two counseling datasets: the High-Low Quality Counseling dataset Pérez-Rosas et al. (2019) and AnnoMI Wu et al. (2022), from which we extract 213 high- and 87 low-quality sessions as specified in the original work. As ClientCAST requires reference dialogues during simulation, we directly use the original counseling sessions from these datasets. Each seeker profile consists of three components: (1) Problems & Reasons for Visiting, (2) Displayed Symptoms from a predefined set of 61 client symptoms, and (3) Apparent Traits, assigned to promote diversity across seeker characteristics.
@@ Appendix I Profile construction for each Seeker Simulator
Seeker Simulator TSNE Hull Area $\uparrow$ PCA Hull Area $\uparrow$
ClientCAST 2954.562 0.312
ESC-Judge 1878.463 0.334
ESC-Role 3093.516 0.378
@@
Appendix L Datasets Used for Fine-tuning Supporter Models
The supporter models were fine-tuned on publicly available emotional support datasets, including ESConv Liu et al. (2021), ExTES Zheng et al. (2024), Psych8k Liu et al. (2023), PsyInsight Chen et al. (2025), and CounselChat Bertagnolli (2020).
Appendix M Human Evaluation Details
Experimental Setup
We selected Ours and ESC-Judge as the two seeker simulators, as they produced the largest performance gaps in supporter evaluation. For supporter models, we included GPT-5-mini and Llama-ESConv, representing prompt-based and fine-tuned supporters, respectively. From the dialogues generated between these seeker simulators and supporter models, we sampled 60 dialogues ensuring a uniform distribution of score discrepancies, covering the full spectrum from cases with similar scores to those with substantial differences. Beyond the $\Delta$-based analysis presented in Section 6.5, we additionally report Human–LLM score correlations and inter-annotator agreement to verify the reliability of the automated evaluation and the human evaluation itself, respectively.
Human–LLM Raw Score Correlations
To assess whether the automated scoring aligns with human judgments at the absolute score level, we computed Pearson and Spearman correlations between human and automated scores. The resulting correlations were moderate to strong across all ES-skills (all $p<0.001$), indicating that the automated evaluation reliably reflects human assessments. (Table 8)
Metric Pearson $r$ Spearman $\rho$
Identification 0.650 0.628
Comforting 0.589 0.546
Suggestions 0.730 0.761
Exp. Sharing 0.636 0.541
Informativeness 0.807 0.794
Table 8: Pearson and Spearman correlations between human and automated evaluation scores.
Inter-Annotator Agreement
To verify the reliability of the human evaluation itself, we assessed inter-annotator agreement using quadratic weighted Cohen’s $\kappa$, which penalizes larger disagreements more heavily. Agreement ranged from fair to strong across metrics, indicating generally reliable human evaluation. (Table 9)
Metric Quadratic $\kappa$
Identification 0.595
Comforting 0.345
Suggestions 0.658
Exp. Sharing 0.821
Informativeness 0.774
Table 9: Inter-annotator agreement measured by quadratic weighted Cohen’s $\kappa$.
@@
Model Coping Utterance Engagement Verbosity Self-Disclosure Resistance Profanity Turns
GPT-4.1-mini 0.33 0.36 0.27 0.28 0.19 0.24 0.58 0.16
Llama-3-8B-Instruct 0.29 0.24 0.27 0.11 0.14 0.29 0.58 0.15
Qwen-2.5-14B-Instruct 0.25 0.32 0.29 0.33 0.22 0.24 0.47 0.15
GPT-5 0.23 0.33 0.27 0.23 0.25 0.25 0.84 0.15
DeepSeek-V3.2 0.39 0.40 0.41 0.52 0.27 0.37 0.91 0.18
SFT 0.38 0.36 0.53 0.66 0.40 0.37 0.66 0.56
Contrastive Learning 0.30 0.32 0.39 0.54 0.32 0.35 0.64 0.83
Ours 0.44 0.47 0.58 0.69 0.45 0.43 0.63 0.74
@@
Llama-3-8B-Instruct ClientCAST 4.527 4.967 4.133 2.463 4.030 4.987 4.997 4.990 4.987 4.990
@@
Llama-CounselChat ClientCAST 4.113 4.563 3.960 2.367 4.010 4.947 4.933 4.767 4.737 4.760
@@
Llama-ESConv ClientCAST 3.900 4.220 3.620 2.383 3.353 4.900 4.880 4.640 4.587 4.653
@@
Llama-ExTES ClientCAST 4.427 4.897 3.993 2.257 3.997 4.993 4.993 4.983 4.973 4.980
@@
Llama-Psych8k ClientCAST 4.237 4.720 3.977 2.143 3.937 4.993 4.993 4.977 4.960 4.973
@@
Llama-PsyInsight ClientCAST 4.070 4.243 3.617 1.743 2.847 4.970 4.967 4.900 4.870 4.900